# TRA: Better Length Generalisation with Threshold Relative Attention

### v1

@@ 2.3 Parameterising D
As each $\delta_{i}$ is a sigmoided scalar value the more times it is multiplied with itself the smaller it gets. The closer $\delta_{i}$ is to 1 the longer the decay takes, the closer to zero the shorter. This provides the model a temporal memory. Enabling it to forget irrelevant past information if necessary, or to treat all timesteps uniformly if temporal ordering is not a concern. We opt for a recency bias for several reasons. Firstly, it has been generally shown to be useful when modeling natural language (Shen et al., 2018; Press et al., 2021; Zhu et al., 2021; Clark et al., 2025). Secondly, it is unclear whether a token occupying the ith position provides useful information beyond causal ordering. The first occurrence can be identified by proximity to the start of sequence token, and middle positions are better differentiated via their unique semantic context as described in Ebrahimi et al. (2024). Finally, decaying memory is the mechanism empowering the resurgence in state space models (Peng et al., 2023; Orvieto et al., 2023; De et al., 2024), recent successful recursive networks (Opper & Siddharth, 2025), and has shown promise with transformers (Csordás et al., 2021; Lin et al., 2025).
@@ 2.4 The Final Attention Weights
$\bm{A^{\prime}}$ $=\text{softmax}\left(\bm{\bar{A}}/\right)$ (11)
This gives us the full TRA mechanism. Note keys which do not meet the threshold are completely masked out from the final attention softmax, and distances are only considered between the query and the keys that matter. This means that TRA exhibits both selective sparsity and allows semantic content to determine and consequently synergise with the positional bias. We provide a Pytorch implementation of TRA in A.
@@ 3.1 Tasks
Flip-Flop Language Modelling: Introduced by Liu et al. (2023a) Flip-Flop language modelling is a algorithmic reasoning benchmark designed to test transformers ability to glitchlessly handle sequential dependencies. Sequences consist of symbol alphabet of three instructions write (w), ignore (i) and read (r) each of which is followed by a bit. To solve the task the model has to recall the bit that follows the nearest write instruction while ignoring all irrelevant information (e.g. in: w1i0i1i1i0i1i0i0i1i0r out: 1). Memorising sequences will lead to failure. The model has to learn a program to succeed. The ability to model a flip-flop is the foundation for all syntactic parsing and algorithmic reasoning capabilities (it necessitates the propagation of information according to abstract rules), and is a pre-requisite to being able to maintain a state hierarchically (Merrill et al., 2022; Liu et al., 2023b). The task consists of three test sets: IID, and two out of distribution sets that vary the number of intervening ignore instructions. OOD sparse increases the number of ignore instructions and requires the ability to handle increased dependency distance. OOD dense decreases the number of ignore instructions and consequently probes whether the model retain focus in the presence of an increased number of attractors (i.e. write instructions). Input examples consist of strings of length 512 with the probability of an ignore instruction being generated set to 0.98 in the sparse set and 0.1 in the dense set. The criterion for success is perfection (i.e. 100% accuracy across all test sets). The authors find that transformers exhibit reasoning errors across a wide range of architectural variants and that issue is scale-invariant, persisting even up to GPT-4.
@@ 3.2 Baselines and Experimental Setup
For both TRA and the baselines we use the same core transformer backbone based on the Llama 3 variant of the architecture (Dubey et al., 2024) consisting on SwiGLU layers Shazeer (2020) coupled with RMSNorm (Zhang & Sennrich, 2019). We compare TRA with the following baselines:
No positional encoding (NoPE): Kazemnejad et al. (2023) claim that decoder only transformers are implicitly able to represent positional information and that therefore explicitly encoding it is unnecessary. Our first baseline is therefore simply standard causal attention with no positional embedding.
@@
Forgetting Transformer (FoT): Utilised by the recent Forgetting Transformer (FoT) (Lin et al., 2025) this mechanism shares the clostest similarity to TRA. The crucial difference is that the forgetting transformer uses standard relative distance compared with TRA contextualised alternative. It therefore serves as the truest test as to whether the issues we hypothesise regarding position are valid.
@@ 5 Conclusion
In this work we investigate whether fundamental issues with length generalisation occur due to a conflict between positional and semantic information. As a probe for this hypothesis we present TRA, a simple modification to the attention mechanism designed to allow both types of information to operate in tandem. We find that the introduction of such a change significantly improves out of distribution generalisation. Moreover, the core innovation of TRA comes from marrying selectively sparsity with contextualised relative distance. We see substantial opportunity for development of future architectures along this axis. Selective sparsity allows columns to focus on what’s important, while the forget gate enables a memory over past time steps. While this latter mechanism does not enable full state tracking behaviour, temporal memory mechanisms have shown enhanced ability to approximate $\text{NC}^{1}$ solutions compared with standard transformers (Merrill et al., 2024). This means that combining a more sophisticated memory with more flexible sparsity selection should yield models which are able to approximate sophisticated algorithmic reasoning more cleanly and with fewer layers. Pursuing research in this direction has the promise of bridging the gap between our theoretical notions of what transformers should be able to do and observed empirical realities, and in so doing yield models which are capable of learning more robust, generalising attentive circuits.

### v2

@@ 2.3 Parameterising D
As each $\delta_{i}$ is a sigmoided scalar value, the more times it is multiplied with itself the smaller it gets. The closer $\delta_{i}$ is to 1 the longer the decay takes, the closer to zero the shorter. This provides the model a temporal memory. Enabling it to forget irrelevant past information if necessary, or to treat all timesteps uniformly if temporal ordering is not a concern. We opt for a recency bias for several reasons. Firstly, it has been generally shown to be useful when modeling natural language (Shen et al., 2018; Press et al., 2021; Zhu et al., 2021; Clark et al., 2025). Secondly, it is unclear whether a token occupying the ith position provides useful information beyond causal ordering. The first occurrence can be identified by proximity to the start of sequence token, and middle positions are better differentiated via their unique semantic context as described in Ebrahimi et al. (2024). Finally, decaying memory is the mechanism empowering the resurgence in state space models (Peng et al., 2023; Orvieto et al., 2023; De et al., 2024), recent successful recursive networks (Opper & Siddharth, 2025), and has shown promise with transformers (Csordás et al., 2021; Lin et al., 2025).
@@ 2.4 The Final Attention Weights
$\bm{A^{\prime}}$ $=\text{softmax}\left(\bm{\bar{A}}\right)$ (11)
This gives us the full TRA mechanism. Note keys which do not meet the threshold are completely masked out from the final attention softmax, and distances are only considered between the query and the keys that matter. This means that TRA exhibits both selective sparsity and allows semantic content to determine and consequently synergise with the positional bias.
@@ 3.1 Tasks
Flip-Flop Language Modelling: Introduced by Liu et al. (2023a) Flip-Flop language modelling is a algorithmic reasoning benchmark designed to test transformers ability to glitchlessly handle sequential dependencies. Sequences consist of symbol alphabet of three instructions: write (w), ignore (i) and read (r) each of which is followed by a bit. To solve the task the model has to recall the bit that follows the nearest write instruction while ignoring all irrelevant information (e.g. in: w1i0i1i1i0i1i0i0i1i0r out: 1). Memorising sequences will lead to failure. The model has to learn a program to succeed. The ability to model a flip-flop is the foundation for all syntactic parsing and algorithmic reasoning capabilities (it necessitates the propagation of information according to abstract rules), and is a pre-requisite to being able to maintain a state hierarchically (Merrill et al., 2022; Liu et al., 2023b). The task consists of three test sets: IID, and two out of distribution sets that vary the number of intervening ignore instructions. OOD sparse increases the number of ignore instructions and requires the ability to handle increased dependency distance. OOD dense decreases the number of ignore instructions and consequently probes whether the model retain focus in the presence of an increased number of attractors (i.e. write instructions). Input examples consist of strings of length 512 with the probability of an ignore instruction being generated set to 0.98 in the sparse set and 0.1 in the dense set. The criterion for success is perfection (i.e. 100% accuracy across all test sets). The authors find that transformers exhibit reasoning errors across a wide range of architectural variants and that the issue is scale-invariant, persisting even up to GPT-4.
@@ 3.2 Baselines and Experimental Setup
For both TRA and the baselines we use the same core transformer backbone based on the Llama 3 variant of the architecture (Dubey et al., 2024); consisting on SwiGLU layers Shazeer (2020) coupled with RMSNorm (Zhang & Sennrich, 2019). We compare TRA with the following baselines:
No positional encoding (NoPE): Kazemnejad et al. (2023) claim that decoder only transformers are implicitly able to represent positional information and that explicitly encoding it is unnecessary. Our first baseline is therefore simply standard causal attention with no positional embedding.
@@
Forgetting Transformer (FoT): The recent Forgetting Transformer (FoT) (Lin et al., 2025) utilises a forget gate as its positional encoding . The crucial difference between it and TRA is that the forgetting transformer uses standard relative distance compared with TRA contextualised alternative. It therefore serves as the truest test as to whether the issues we hypothesise regarding position are valid.
@@ 5 Conclusion
In this work we investigate whether fundamental issues with length generalisation occur due to a conflict between positional and semantic information. As a probe for this hypothesis we present TRA, a simple modification to the attention mechanism designed to allow both types of information to operate in tandem. We find that the introduction of such a change significantly improves out of distribution generalisation. Moreover, the core innovation of TRA comes from marrying selectively sparsity with contextualised relative distance. We see substantial opportunity for development of future architectures along this axis. Selective sparsity allows columns to focus on what’s important, while the forget gate enables a memory over past time steps. While this latter mechanism does not enable full state tracking behaviour, temporal memory mechanisms have shown enhanced ability to approximate $\text{NC}^{1}$ solutions compared with standard transformers (Merrill et al., 2024). This means that combining a more sophisticated memory with a more flexible sparsification threshold should yield models which are able to approximate complex algorithmic reasoning more cleanly and with fewer layers. Pursuing research in this direction has the promise of bridging the gap between our theoretical notions of what transformers should be able to do and observed empirical realities, and in so doing yield models which are capable of learning more robust, generalising attentive circuits.